If consumption is C=100+0.75Yd




Taxes is T=50+0.5Y




Export is X=200




Import is M=50+0.25Y




Government spending is G=150




Investment is I=200

.Use the multiplier applicable to export,to explain how a100–billion decline in demand


for export could affect the economy’s:


(i)


GDP/income




1
Expert's answer
2021-05-16T19:18:10-0400


An export multiplier is a multiplier that is applicable to export. The effect of a decline in demand by 100 billion on GDP/income will be computed by the export multiplier. Thus,


"Export \space Multiplier=\frac{1}{1-MPC+M}"


M is the marginal propensity to import which is calculated by dividing the change in spending on imports by change in income. According to the question, the change in spending on imports is 0.25 and the change in income is 0.5. Thus,


"M=\frac{change\space in\space import spending}{change\space in\space income}"


"=\frac{0.25}{0.5}"


"=0.5"


 Hence, the marginal propensity to import is 0.5.


Therefore the export multiplier will be:


"=\frac{1}{1-0.75+0.5}\\=1.33"


The export multiplier shows that the GDP/income will be decreased by 1.33 billion with the decline in the 100-billion decline in the demand.